OpenStudy (anonymous):

M is the midpoint of AB. Find the coordinates of B given A(2,-1) and M(8,3).

OpenStudy (anonymous):

Use the mid-point formula: \[M=(\frac{ {x_{1}+x_{2}} }{ 2 },\frac{ y_{1}+y_{2} }{ 2 })\] Plug in what you have, which is Midpoint(8,3) and A(2,-1) Solve for missing parts, which is point B \[8=\frac{ 2+x }{ 2}, 3=\frac{ -1+y }{ 2 }\] \[(2)8=(\frac{ 2+x }{ 2})2, (2)3=(\frac{ -1+y }{ 2 })2\] \[16=2+x, 6=-1+y\] -2 -2 , +1 +1 x = 18 , y = 7 B(18,7)

OpenStudy (anonymous):

Thank you very much i just needed the midpoint formula thanks now i can check my work. :)

OpenStudy (anonymous):

Yup, good luck!